Assume a certain stock is trading above the strike price on a call option and is...

  1. Assume a certain stock is trading above the strike price on a call option and is assumed to have 0% volatility in the future. Also assume interest rates are 0%. Write a simplified equation to find the value of that call option.
Answer #1

d1=(ln(S/K)+(r+vol^2/2)*T)/(vol*sqrt(T))=infinite
Hence, N(d1)=1

d2=(ln(S/K)+(r-vol^2/2)*T)/(vol*sqrt(T))=infinite
Hence, N(d2)=1

Price of call option=SN(d1)-Ke^(-rt)*N(d2)
=SN(d1)-KN(d2)
=S-K

where S is the spot price at the time of valuation and K is the strike price